OpenStudy (anonymous):

how would I solve this? x = 5 sin t, y = 5 cos t

hartnn (hartnn):

you want to eliminate 't' ? just square and add them

OpenStudy (gorv):

square them and add

OpenStudy (gorv):

sin^2t+cos^2t=1 trigonometry formula

OpenStudy (anonymous):

so i sq them both?

OpenStudy (anonymous):

wait what?

hartnn (hartnn):

yes find x^2 and y^2

hartnn (hartnn):

\(x =5 \sin t \\ x^2 = (5 \sin t)^2 = ... ?\)

OpenStudy (anonymous):

ok 25 sin^2 but how do i add them

hartnn (hartnn):

y^2 = 25 cos^2 t now add them \( x^2 +y^2 = 25 \sin^2 t +25 \cos^2 t\) what can you factor out ??

OpenStudy (anonymous):

25 and t

hartnn (hartnn):

note that sin t and cos t are functions and are not separate entities like 'sin' and 't' so you cannot factor out t :)

hartnn (hartnn):

just factor out 25 and see what u get

hartnn (hartnn):

and use the fact that \( \sin^2 t + \cos^2 t =1\)

OpenStudy (anonymous):

so it would be a circle

hartnn (hartnn):

yes

hartnn (hartnn):

what would be the radius of that circle ?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

but what does this mean: Eliminate the parameter t from the following.

OpenStudy (phi):

It means to do what you just did. you now have x^2 + y^2 = 5^2 no t !

OpenStudy (anonymous):

no cos or sin?

hartnn (hartnn):

no 't'

hartnn (hartnn):

eliminate 't' means combine the equations in such a way that the new equation does not have the variable 't' at all

OpenStudy (anonymous):

so the answer looks like this 5sin^2 +5cos^2

hartnn (hartnn):

no

OpenStudy (anonymous):

geez i dont get it

hartnn (hartnn):

\(\sin^2 t +\cos^2 t\) is already 1

OpenStudy (anonymous):

oh yeah

hartnn (hartnn):

25 (sin^2 t + cos^2 t) = 25*1 =25

OpenStudy (anonymous):

25 (sin^2 t + cos^2 t) = 25

OpenStudy (anonymous):

thats what they are asking for?

OpenStudy (phi):

I think they are asking for you to change x = 5 sin t, y = 5 cos t into x^2 + y^2 = 25

hartnn (hartnn):

^^

OpenStudy (phi):

the first form x = 5 sin t, y = 5 cos t is "parametric form" you can calculate (x,y) pairs for values of "t" the other form x^2 + y^2 = 25 has no "t". You can still find (x,y) pairs, but it is a bit more painful

OpenStudy (anonymous):

thank you so much thanks for your brilliance

OpenStudy (phi):

If you want more background, see https://www.khanacademy.org/math/precalculus/parametric_equations and look at the first video in each section.

OpenStudy (anonymous):

will do

OpenStudy (phi):

and the 3rd video in the first section looks close to your problem https://www.khanacademy.org/math/precalculus/parametric_equations/parametric/v/parametric-equations-3 The first 2 videos talk about the parametric form... which is useful if the idea is new.

OpenStudy (anonymous):

well i am stuck on a another problem if you guys are up for the challenge

OpenStudy (anonymous):

Find all solutions if 0 ≤ x < 2π. Use exact values only. (Enter your answers as a comma-separated list.) sin x cos 2x + cos x sin 2x = 1/2